OpenStudy (anonymous):

Choose the Standard Form of the equation Y= −1/3x - 4.

OpenStudy (anonymous):

options please!!!!!!

OpenStudy (lgbasallote):

multiply every term by 3

OpenStudy (anonymous):

the std is y=mx+h or y-y0=m(x-x0)?

OpenStudy (lgbasallote):

y = mx + h is slope intercept form standard from is ax + by + c =0